The maximum velocity at which a truck can safely travel without toppling over on a curve of radius 250 m ( the height of the centre of gravity of the truck above the ground is 1.5 m and the distance between the wheels is 1.5 m the track being horizontal) is :

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Net moment  about centre of gravity=>
 m * 0.75 = m * ( v ² / r ) * 1.5
we get,
v = root ( ( r * g * 0.75 ) / 1.5 ) 
Putting all the values we get 
v = 35.0 m/s